6. Durran has recently acquired a rare art piece that he plans to put on display in his private collection. He estimates that revenues generated from donations and admissions tickets to see the new exhibit will be $2,000 per month for the next eight years. If he requires a rate of return of 12%, how much are the expected cash flows worth for him today? Round to two decimal places. Adelai Inc. loans you S32,000 for five years to buy a car. The loan must be repaid in 60 equal monthly payments. The annual interest rate on the loan is 8percent. What is your monthly payment? Round to two decimal places. 7. Your company has received a $50,000 loan from an industrial finance company. The annual payments are S6,202.70. If the company is paying 9 percent interest per year, how many loan payments must the company make? Round to the nearest number of periods. 8. You are ready to retire. A glance at your 401 (k) statement indicates that you have $2,750,000. If the funds remain in an account earning 4.5% annually, how much could you withdraw at the end of each year for the next 30 years? Round to two decimal places. 9. 10. If you wish to accumulate $280,000 in the childs college fund after 18 years, and can invest at a 7.5% annual rate, how much must you invest at the end of each year if the first deposit is made at the end of the first year? Round to two decimal places.
Answer #1

Answer:

Question 6)

Given

Monthly revenue R=$2000

Number of years =8 years

So n=8*12=96 months

Interest rate=12% per annum

r=12%/12=1%

So PV of cash flow= R*(1-(1+r)^-n)/r=2000*(1-(1+1%)^-96)/1%=$123,055.41